OpenStudy (anonymous):

log(x-4) - logx = 2log3

OpenStudy (anonymous):

I got -1/2 but that's not right

OpenStudy (anonymous):

\[log(x-4)-log(x)=2log(3)\] \[log(\frac{x-4}{x})=\log(3^2)\]

myininaya (myininaya):

\[\frac{x-4}{x}=3^2\]

myininaya (myininaya):

\[x-4=9x\]

OpenStudy (anonymous):

\[\frac{x-4}{x}=9\] \[x-4=9x\] \[8x=-4\] \[x=-\frac{1}{2}\] which certainly isn't right because you cannot take the log of a negative number

myininaya (myininaya):

no solution!

myininaya (myininaya):

or say no real solution

OpenStudy (anonymous):

\[\color{blue}{\text{ so i guess there is no solution}}\]

myininaya (myininaya):

no so you know there is no real solution

OpenStudy (anonymous):

\[\color{#FF0088}{\text{ in this color}}\]

OpenStudy (anonymous):

okay that's what I thought, thanks XD